# Bondi mass, memory effect and balance law of polyhomogeneous spacetime

### v2

@@ 4. Bondi mass of smooth asympotically spacetime
We have reexamined the Bondi mass at null infinity via the Iyer-Wald formula within the Bondi-Sachs coordinates.
5. Bondi mass of polyhomogeneous spacetimes
For smooth asymptotically flat spacetimes, we have successfully employed the Iyer-Wald framework to obtain the charge associated with the asymptotic time translation at null infinity in the previous section. In 1995, Chruściel, MacCallum and Singleton discussed the Bondi mass of a class of polyhomogeneous spacetimes with $N_{3}=-\infty$ [13]. Recently, the BMS charges in special polyhomogeneous spacetimes are discussed by using the Barnich-Brandt prescription [31]. In this section, we attempt to derive the mass expression of polyhomogeneous spacetimes by using the Iyer-Wald formula.
Recall that the Iyer-Wald canonical energy is given by [19, Eq. (83)]
@@ 5. Bondi mass of polyhomogeneous spacetimes
Comparing with the original definition in [19], there appears a total minus in the expression of the canonical energy. This is because we take the most minus signature.
@@
$\begin{split}-\int_{\infty}Q[\xi]&=\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}\nabla^{c}\xi^{d}\\ &=\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\nabla_{e}(\frac{\partial}{\partial u})^{d}\\ &=\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\Gamma^{d}_{\ ef}(\frac{\partial}{\partial u})^{f}\\ &=\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\Gamma^{d}_{\ e0}\\ &=\frac{1}{16\pi}\int_{\infty}\epsilon_{cdab}g^{ce}\Gamma^{d}_{\ e0}\\ &=\frac{1}{16\pi}\int_{\infty}\big{(}\epsilon_{01ab}g^{0e}\Gamma^{1}_{\ e0}+\epsilon_{10ab}g^{1e}\Gamma^{0}_{\ e0}\big{)}\\ &=\frac{1}{16\pi}\int_{\infty}\sqrt{-g}\big{(}g^{0e}\Gamma^{1}_{\ e0}-g^{1e}\Gamma^{0}_{\ e0}\big{)}(\mathrm{d}\theta\wedge\mathrm{d}\varphi)_{ab}.\end{split}$ (5.3)
@@
$-Q[\xi]=\frac{1}{8\pi}\lim_{r\rightarrow\infty}\int_{S^{2}}\big{[}U_{1}-U_{1}^{\prime}\big{]}\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.$ (5.6)
@@
$-Q[\xi]=\frac{1}{8\pi}\lim_{r\rightarrow\infty}\int_{S^{2}}U_{1,0}\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.$ (5.7)
@@
$\begin{split}\int_{\infty}\xi\cdot\Theta&=\int_{\infty}(\frac{\partial}{\partial u})^{a}\Theta_{abc}\\ &=\frac{1}{16\pi}\int_{\infty}(\frac{\partial}{\partial u})^{a}\epsilon_{dabc}g^{de}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\epsilon_{0dbc}g^{de}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{1e}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{1e}g^{fh}\bigg{(}\partial_{f}\delta g_{eh}-\Gamma^{a}_{\ fe}\delta g_{ah}-\Gamma^{a}_{\ fh}\delta g_{ea}\\ &\hskip 56.9055pt-\partial_{e}\delta g_{fh}+\Gamma^{a}_{\ ef}\delta g_{ah}+\Gamma^{a}_{\ eh}\delta g_{fa}\bigg{)}\\ &=-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{1e}g^{fh}\bigg{(}\partial_{f}\delta g_{eh}-\Gamma^{a}_{\ fh}\delta g_{ea}-\partial_{e}\delta g_{fh}+\Gamma^{a}_{\ eh}\delta g_{fa}\bigg{)}\\ \end{split}$ (5.8)
@@
$\begin{split}&\bullet\ \delta g_{uu}\sim O\big{(}r^{-1}(\ln r)^{N_{3}+1}\big{)},\\ &\bullet\ \delta g_{ru}=\delta g_{rr}=\delta g_{r\theta}=\delta g_{r\varphi}=0,\\ &\bullet\ \delta g_{u\theta},\delta g_{u\varphi}\sim O\big{(}(\ln r)^{N_{3}+1}\big{)},\\ &\bullet\ \delta g_{\theta\theta},\delta g_{\theta\varphi},\delta g_{\varphi\varphi}\sim O\big{(}r(\ln r)^{N_{3}+1}\big{)}.\end{split}$ (5.10)
@@
$\mathcal{E}=-\frac{1}{8\pi}\int_{S^{2}}(\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0})\mathrm{d}S.$ (5.20)
This is just the ‘mass’ introduced by Newman and Unti in [6]. Moreover, recall that
@@
$\begin{split}\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}u}&=\frac{1}{8\pi}\int_{S^{2}}\bigg{(}\eth^{2}\dot{\bar{\sigma}}_{2,0}+\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}+\bar{\eth}^{2}\dot{\sigma}_{2,0}+\bar{\sigma}_{2,0}\ddot{\sigma}_{2,0}\bigg{)}\mathrm{d}S\\ &=\frac{1}{8\pi}\int_{S^{2}}\bigg{(}\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\ddot{\sigma}_{2,0}\bigg{)}\mathrm{d}S\\ &=\frac{1}{8\pi}\int_{S^{2}}\bigg{[}\frac{\mathrm{d}}{\mathrm{d}u}(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0})-2|\dot{\sigma}_{2,0}|^{2}\bigg{]}\mathrm{d}S.\\ \end{split}$ (5.22)
@@
^$\begin{split}\hat{\mathcal{E}}&\equiv\mathcal{E}-\frac{1}{8\pi}\int_{S^{2}}\bigg{(}\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0}\bigg{)}\mathrm{d}S\\ &=-\frac{1}{8\pi}\int_{S^{2}}\bigg{(}\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0}+\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0}\bigg{)}\mathrm{d}S.\end{split}$ (5.23)
@@
$\frac{\mathrm{d}\hat{\mathcal{E}}}{\mathrm{d}u}=-\frac{1}{4\pi}\int_{S^{2}}|\dot{\sigma}_{2,0}|^{2}\ \mathrm{d}S\leq 0.$ (5.24)
In fact, $\hat{\mathcal{E}}$ is just the well-known Bondi mass.
@@
$\begin{split}\mathcal{E}=&\frac{1}{4\pi}\int_{S^{2}}\bigg{[}\frac{3}{4}(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3})-\frac{1}{2}(\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0})\bigg{]}\mathrm{d}S\\ &+\frac{1}{8\pi}\int_{S^{2}}(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3})z\ \mathrm{d}S.\end{split}$ (5.31)
@@
$\begin{split}\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}u}&=\frac{1}{4\pi}\int_{S^{2}}\bigg{[}\frac{3}{4}(\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\bar{\sigma}}_{2,0}\dot{\Psi}_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}+\dot{\sigma}_{2,0}\dot{\bar{\Psi}}_{0}^{3,0})\\ &\ \ \ \ \ \ \ \ \ \ \ \ -\frac{1}{2}(\dot{\Psi}_{2}^{3,0}+\dot{\bar{\Psi}}_{2}^{3,0})\bigg{]}\mathrm{d}S\\ &\ \ +\frac{1}{8\pi}\int_{S^{2}}\bigg{[}\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\bar{\sigma}}_{2,0}\dot{\Psi}_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}+\dot{\sigma}_{2,0}\dot{\bar{\Psi}}_{0}^{3,0}\bigg{]}z\ \mathrm{d}S.\end{split}$ (5.33)
@@
$\begin{split}\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}u}&=\frac{1}{4\pi}\int_{S^{2}}\bigg{[}\frac{3}{4}(\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+\frac{1}{2}(\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\ddot{\sigma}_{2,0})\bigg{]}\mathrm{d}S\\ &\ \ +\frac{1}{8\pi}\int_{S^{2}}\bigg{[}\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg{]}z\ \mathrm{d}S\\ &=\frac{1}{16\pi}\int_{S^{2}}\bigg{[}\partial_{u}\bigg{(}3(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+2(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0})\bigg{)}\\ &\ \ \ \ \ -4|\dot{\sigma}_{2,0}|^{2}\bigg{]}\mathrm{d}S+\frac{1}{8\pi}\int_{S^{2}}\partial_{u}\bigg{(}\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg{)}z\ \mathrm{d}S.\end{split}$ (5.35)
@@
^$\begin{split}\hat{\mathcal{E}}&\equiv\mathcal{E}-\frac{1}{16\pi}\int_{S^{2}}\bigg{(}3(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+2(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0})\bigg{)}\mathrm{d}S\\ &\ \ \ \ \ \ -\frac{1}{8\pi}\int_{S^{2}}\bigg{(}\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg{)}z\ \mathrm{d}S.\end{split}$ (5.36)
@@
$\frac{\mathrm{d}\hat{\mathcal{E}}}{\mathrm{d}u}=-\frac{1}{4\pi}\int_{S^{2}}|\dot{\sigma}_{2,0}|^{2}\ \mathrm{d}S\leq 0.$ (5.37)
6. Gravitational wave memory and balance law
The linear memory of gravitational waves (GW) arising from the change in the quadrupole moment of a gravitational wave source was identified by Zel’dovich et al. [22, 32, 33, 34]. Subsequently, Christodoulou and Frauendiener discovered that gravitational waves themselves can generate memory known as the nonlinear memory [23, 24]. The effect of this kind of memory is related to the relative displacement of nearby observers, so it is also called the displacement memory. Over the past few years, numerous works have been established on the potential for detecting the nonlinear memory [35, 36, 37, 38, 39, 40, 41, 42]. However, gravitational wave memory measured by the detector is notably weak, as GW memory behaves mainly as a quasi-direct current signal [28]. Consequently, the development of an accurate theoretical model is crucial for the detection of GW memory. Noticing that the BMS theory does not necessitate the slow and weak field conditions for the GW sources, Nichols et al. proposed a method to calculate the GW memory based on the BMS theory [25, 26]. More recently, combining the results of numerical relativity, Cao et al. applied the BMS method to develop a surrogate model that correlates the parameters of binary black holes with GW memory [28]. Utilizing this surrogate model, they estimated the GW memory associated with all 48 binary black hole events documented in GWTC-2. In the context of the BMS method for calculating the GW memory, a critical step involves the derivation of the balance law of smooth asymptotically flat spacetimes. For polyhomogeneous spacetimes, it is worthwhile reexamining the balance law, as the asymptotic behaviors of Newman-Penrose (NP) quantities are significantly different between polyhomogeneous spacetimes and smooth asymptotically flat spacetimes.

### v3

@@ 4 Bondi mass of smooth asympotically flat spacetimes
In this section, we have reexamined the Bondi mass at null infinity via the Iyer-Wald formula within the Bondi-Sachs coordinates. Our analysis demonstrates that the Iyer-Wald formula naturally leads to the Bondi mass expression directly in the physical spacetime, thereby bypassing the need of conformal compactification.
5 Bondi mass of polyhomogeneous spacetimes
For smooth asymptotically flat spacetimes, we have successfully employed the Iyer-Wald framework to obtain the charge associated with the asymptotic time translation at null infinity in the previous section. In 1995, Chruściel, MacCallum and Singleton discussed the Bondi mass of a class of polyhomogeneous spacetimes with $N_{3}=$-$\infty$ Chrusciel95 . Recently, the BMS charges in special polyhomogeneous spacetimes are discussed by using the Barnich-Brandt prescription Godazagar2020 . In this section, we attempt to derive the mass expression of polyhomogeneous spacetimes by using the Iyer-Wald formula.
Recall that the Iyer-Wald canonical energy is given by (Iyer94, , Eq. (83))
@@ 5 Bondi mass of polyhomogeneous spacetimes
Comparing with the original definition in Iyer94 , there appears a total minus in the expression of the canonical energy. This is because we take the Landau-Lifshitz timelike convention for the signature of the spacetime metric, (+,,,)(+,-,-,-).
@@
$\begin{split}-\int_{\infty}Q[\xi]&=\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}\nabla^{c}\xi^{d}\\ &=\frac{1}{16\pi}\int_{\infty}\sqrt{-g}\big(g^{0e}\Gamma^{1}_{\ e0}-g^{1e}\Gamma^{0}_{\ e0}\big)(\mathrm{d}\theta\wedge\mathrm{d}\varphi)_{ab}.\end{split}$ (5.3)
@@
$Q[\xi]=\frac{1}{8\pi}\lim_{r\rightarrow\infty}\int_{S^{2}}\big[U_{1}-U_{1}^{\prime}\big]\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.$ (5.6)
@@
$Q[\xi]=\frac{1}{8\pi}\lim_{r\rightarrow\infty}\int_{S^{2}}U_{1,0}\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.$ (5.7)
@@
[…]
@@
$\begin{split}&\delta g_{uu}\sim O\big(r^{-1}(\ln r)^{N_{3}+1}\big),\\ &\delta g_{ru}=\delta g_{rr}=\delta g_{r\theta}=\delta g_{r\varphi}=0,\\ &\delta g_{u\theta},\delta g_{u\varphi}\sim O\big((\ln r)^{N_{3}+1}\big),\\ &\delta g_{\theta\theta},\delta g_{\theta\varphi},\delta g_{\varphi\varphi}\sim O\big(r(\ln r)^{N_{3}+1}\big).\end{split}$ (5.10)
@@
$\mathcal{E}=-\frac{1}{8\pi}\int_{S^{2}}(\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0})\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.$ (5.20)
This is just the ‘mass’ introduced by Newman and Unti in NU62 . Moreover, recall that
@@
[…]
@@
$\begin{split}\hat{\mathcal{E}}&\equiv\mathcal{E}-\frac{1}{8\pi}\int_{S^{2}}\bigg(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0}\bigg)\sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &=-\frac{1}{8\pi}\int_{S^{2}}\bigg(\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0}+\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0}\bigg)\sin\theta\mathrm{d}\theta\mathrm{d}\varphi.\end{split}$ (5.23)
@@
$\frac{\mathrm{d}\hat{\mathcal{E}}}{\mathrm{d}u}=-\frac{1}{4\pi}\int_{S^{2}}|\dot{\sigma}_{2,0}|^{2}\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi\leq 0.$ (5.24)
In fact, $\hat{\mathcal{E}}$ is just the well-known Bondi mass. It turns out that this charge coincides with the one obtained by Godazgar and Macaulay in (Godazagar2020, , Eq.(5.10)), where the BMS charges for polyhomogeneous spacetimes with $\Psi_{0}^{3}=0$ are derived by using the Barnich-Brandt prescription Barnich2002 .
@@
$\begin{split}\mathcal{E}=&\frac{1}{4\pi}\int_{S^{2}}\bigg[\frac{3}{4}(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3})-\frac{1}{2}(\Psi_{2}^{3,0}+\bar{\Psi}_{2}^{3,0})\bigg]\sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &+\frac{1}{8\pi}\int_{S^{2}}(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3})z\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi.\end{split}$ (5.31)
@@
$\begin{split}\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}u}&=\frac{1}{4\pi}\int_{S^{2}}\bigg[\frac{3}{4}(\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\bar{\sigma}}_{2,0}\dot{\Psi}_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}+\dot{\sigma}_{2,0}\dot{\bar{\Psi}}_{0}^{3,0})\\ &\ \ \ \ \ \ \ \ \ \ \ \ -\frac{1}{2}(\dot{\Psi}_{2}^{3,0}+\dot{\bar{\Psi}}_{2}^{3,0})\bigg]\sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &\ \ +\frac{1}{8\pi}\int_{S^{2}}\bigg[\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\bar{\sigma}}_{2,0}\dot{\Psi}_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}+\dot{\sigma}_{2,0}\dot{\bar{\Psi}}_{0}^{3,0}\bigg]z\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi.\end{split}$ (5.33)
@@
$\begin{split}\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}u}&=\frac{1}{4\pi}\int_{S^{2}}\bigg[\frac{3}{4}(\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+\frac{1}{2}(\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\ddot{\sigma}_{2,0})\bigg]\sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &\ \ +\frac{1}{8\pi}\int_{S^{2}}\bigg[\ddot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\ddot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg]z\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &=\frac{1}{16\pi}\int_{S^{2}}\bigg[\partial_{u}\bigg(3(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+2(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0})\bigg)\\ &\ \ \ \ \ -4|\dot{\sigma}_{2,0}|^{2}\bigg]\sin\theta\mathrm{d}\theta\mathrm{d}\varphi+\frac{1}{8\pi}\int_{S^{2}}\partial_{u}\bigg(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg)z\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi.\end{split}$ (5.35)
@@
$\begin{split}\hat{\mathcal{E}}&\equiv\mathcal{E}-\frac{1}{16\pi}\int_{S^{2}}\bigg(3(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0})+2(\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}+\bar{\sigma}_{2,0}\dot{\sigma}_{2,0})\bigg)\sin\theta\mathrm{d}\theta\mathrm{d}\varphi\\ &\ \ \ \ \ \ -\frac{1}{8\pi}\int_{S^{2}}\bigg(\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0}+\dot{\sigma}_{2,0}\bar{\Psi}_{0}^{3,0}\bigg)z\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi.\end{split}$ (5.36)
@@
$\frac{\mathrm{d}\hat{\mathcal{E}}}{\mathrm{d}u}=-\frac{1}{4\pi}\int_{S^{2}}|\dot{\sigma}_{2,0}|^{2}\ \sin\theta\mathrm{d}\theta\mathrm{d}\varphi\leq 0.$ (5.37)
In summary, we have obtained the explicit expression for the canonical energy, given in (5.18), for vacuum polyhomogeneous spacetimes by applying the Iyer-Wald formalism within Newman-Unti coordinates. When $\Psi_{0}^{3}=0$, regardless of whether $\Psi_{0}^{4}$ vanishes, this canonical energy reduces to the Newman-Unti mass. By subtracting a suitable quantity, we find that the resulting ‘reduced energy’ (5.23) is just the well-known Bondi mass, which decreases monotonically in the presence of gravitational radiation. Furthermore, for the case $N_{3}=0$, we define the ‘reduced energy’ by (5.36), which also possesses a monotonicity property, as shown in (5.37).
6 Gravitational wave memory and balance law
The linear memory of gravitational waves (GW) arising from the change in the quadrupole moment of a gravitational wave source was identified by Zel’dovich et al. Zel74 ; Payne1983 ; BG1985 ; BT1987 . Subsequently, Christodoulou and Frauendiener discovered that gravitational waves themselves can generate memory known as the nonlinear memory Christo91 ; Frau92 . The effect of this kind of memory is related to the relative displacement of nearby observers, so it is also called the displacement memory. Over the past few years, numerous works have been established on the potential for detecting the nonlinear memory Seto2009 ; PBP2010 ; HL2010 ; CJ2012 ; MCC2014 ; LTLBC2016 ; MTL2017 ; DTLW2020 . However, gravitational wave memory measured by the detector is notably weak, as GW memory behaves mainly as a quasi-direct current signal Cao21-2 . Consequently, the development of an accurate theoretical model is crucial for the detection of GW memory. Noticing that the BMS theory does not necessitate the slow and weak field conditions for the GW sources, Nichols et al. proposed a method to calculate the GW memory based on the BMS theory Nichols17 ; Khera21 . More recently, combining the results of numerical relativity, Cao et al. applied the BMS method to develop a surrogate model that correlates the parameters of binary black holes with GW memory Cao21-2 . Utilizing this surrogate model, they estimated the GW memory associated with all 48 binary black hole events documented in GWTC-2. In the context of the BMS method for calculating the GW memory, a critical step involves the derivation of the balance law of smooth asymptotically flat spacetimes. For polyhomogeneous spacetimes, it is worthwhile reexamining the balance law, as the asymptotic behaviors of Newman-Penrose (NP) quantities are significantly different between polyhomogeneous spacetimes and smooth asymptotically flat spacetimes.